Fing the equation, in slope-intercept form, of a line passing through (4,-3) and is perpendicular to the graph of -1/5x+y=7.
Answer by mananth(16946)

-0.2 x + 1 y = 7
Find the slope of this line

1 y = 0.2 x + 7
Divide by 1
y = 0.200 x + 7
Compare this equation with y=mx+b
slope m = 0.20

The slope of a line perpendicular to the above line will be the negative reciprocal -5
m1*m2=-1
The slope of the required line will be -5.00

m= -5 ,point ( 4 , -3 )
Find b by plugging the values of m & the point in
y=mx+b
-3 = -20.00 + b
b= 17
m= -5
The required equation is y = -5 x + 17
m.ananth@hotmail.ca